If the diameter of a carbon atom is 0.15 nm, calculate the number of carbon atoms which can be placed side by side in a straight line across length of scale of length 20 cm long.

Grade:12

2 Answers

Gaurav
9 years ago
Hello Student
Diameter of C atom = 0.15 x 10-9 m
One carbon atom occupie0.15 x 10-9m
Number of C atoms that can be placed in a straight line
=20 x 10-2 m/0.15 x 10-9m =1.33 x 109
Ramnath Venkatesh
5 years ago
Length of scale = 20 cm=20 x 10-2
Diameter of Carbon atom(given)= 1.5nm=0.15 x 10-9m
No.of atoms to be placed in the line=(20 x 10-2)m/(0.15 x 10-9)m=1.33×109